OpenStudy (anonymous):

determine the value of a so that the average rate of change of the function j(x)= 3x²+ax-4 over the interval -1≤x≤2 is 8

OpenStudy (anonymous):

average rate of change, as in \(\frac{f(2)-f(-1)}{3}=8\) ?

OpenStudy (anonymous):

we can do this for sure

OpenStudy (anonymous):

you want \[\frac{f(2)-f(-1)}{3}=8\] or \[f(2)-f(-1)=24\] so lets compute the left hand side okay?

OpenStudy (anonymous):

\[f(2)=12+2a-4=2a+8\] and \[f(-1)=3-a-4=-a-1\] you with me so far?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so you want \[2a+8-(-a-1)=24\] i.e. \[3a+9=24\]

OpenStudy (anonymous):

oh! and then solve for a

OpenStudy (anonymous):

yeah should take only two steps right?

OpenStudy (anonymous):

3a=24-9 3a=15 a=5

OpenStudy (anonymous):

YAY! awesome thanks ! :)

OpenStudy (anonymous):

that ought to do it, yes

OpenStudy (anonymous):

yw